### v1

# Aligned Probing: Relating Toxic Behavior and Model Internals

Abstract
Warning: This paper contains offensive text.
We introduce aligned probing, a novel interpretability framework that aligns the behavior of language models (LMs), based on their outputs, and their internal representations (internals). Using this framework, we examine over 20 OLMo, Llama, and Mistral models, bridging behavioral and internal perspectives for toxicity for the first time. Our results show that LMs strongly encode information about the toxicity level of inputs and subsequent outputs, particularly in lower layers. Focusing on how unique LMs differ offers both correlative and causal evidence that they generate less toxic output when strongly encoding information about the input toxicity. We also highlight the heterogeneity of toxicity, as model behavior and internals vary across unique attributes such as Threat. Finally, four case studies analyzing detoxification, multi-prompt evaluations, model quantization, and pre-training dynamics underline the practical impact of aligned probing with further concrete insights. Our findings contribute to a more holistic understanding of LMs, both within and beyond the context of toxicity.
1 Introduction
Language models (LMs) may produce toxic text that contains hate speech, insults, or vulgarity, even when prompted with innocuous text (Gehman et al., 2020; de Wynter et al., 2024). Preventing the generation of such toxic language is an important part of making LMs safer to use (Kumar et al., 2023). Efforts in this direction include analyzing the toxicity of model generations (Ousidhoum et al., 2021; Hartvigsen et al., 2022), the effects of pre-training data (Groeneveld et al., 2024; Longpre et al., 2024), and model detoxification (Lee et al., 2024; Li et al., 2024; Yang et al., 2024). However, the scope of such work is limited as they mostly focus on the behavior (Chang and Bergen, 2024) of models based on their outputs, ignoring the model-internal perspective (Hu and Levy, 2023; Waldis et al., 2024b; Mosbach et al., 2024), and they treat toxic language as homogeneous rather than diverse (Pachinger et al., 2023; Wen et al., 2023). Thus, we lack a methodological framework to answer the question:
How do LMs encode information about toxicity, and what is the interplay between their internals and behavior?
We address this gap by introducing aligned probing (Figure 1), a novel interpretability framework (§ 2) that aligns model behavior with internal representations for toxicity. First, we prompt LMs with inputs and assess the toxicity of their generated outputs. Then, during the forward pass, we extract internal representations at each layer to analyze how models encode toxic language. Specifically, we use linear probing (Tenney et al., 2019a; Belinkov, 2022) to train linear models to use these internals to predict specific properties (like input toxicity). Since probes have limited capacity and are rigorously validated (Hewitt and Liang, 2019; Voita and Titov, 2020), their prediction performance on held-out data estimates information strength. Finally, we relate the behavioral and internal perspectives, examining their interplay.
To account for the heterogeneity of toxic language, we consider six fine-grained attributes (§ 3) and show their varying dependence on specific words. For example, Threats rely on context, while Sexually Explicit toxicity is focused on individual words. Using aligned probing and the RealToxicPrompts dataset (Gehman et al., 2020), we evaluate 20+ popular pre-trained and instruction-tuned LMs, including Llama, OLMo, and Mistral. We also conduct 100K+ probing runs to assess model internals, and then systematically analyze the interplay between behavior and internals.
We first examine high-level insights across LMs (§ 4), and show that LMs strongly encode information about the toxicity of text in lower layers. This provides an alternative perspective to previous findings that localize toxicity in upper layers (Lee et al., 2024). We also find that LMs replicate and amplify toxicity more than humans as they strongly encode input toxicity, especially when focused on single words like Profanities.
Next, we analyze individual LMs in detail (§ 5) and find that less toxic models encode more information about input toxicity. We further establish that this is a causal relationship (§ 6), showing that LMs are generally less toxic when they know more about the toxicity of a given input. Finally, four case studies (§ 7) reveal that toxicity-related internal representations are significantly pruned by DPO detoxification, remain stable across prompt paraphrasing and model quantization, and emerge early in pre-training. Our work thus makes the following methodological and empirical contributions to toxicity and interpretability research:
1.
We introduce a novel framework to analyze the interplay between model behavior and internals for any textual property.
2.
We comprehensively study toxicity with 20 contemporary LMs.
3.
We provide in-depth practical insights by comparing different LMs, multi-prompt evaluations, pre-training dynamics, detoxification via DPO, and model quantization.
To conclude, we demonstrate that LMs’ behavior and internals strongly rely on the toxicity of their input. Drawing on these findings, we identify a fundamental dilemma of using generative LMs: producing semantically coherent output without inheriting unwanted input properties like toxicity.
2 Aligned Probing
We introduce aligned probing, an interpretability framework that explicitly aligns model behavior and internals to examine their interplay in the context of toxic language. We first evaluate the behavior of LMs (§ 2.1), based on the toxicity scores ( and ) of the input () and the corresponding output (). Next, we analyze how strongly LMs encode information about these toxicity scores within their internal representations of the input () and output (), extracted during generation (§ 2.2). Finally, we correlate the resulting information strength () with model behavior to investigate their interplay (§ 2.3). While this study focuses on toxic text, language that likely makes people leave a discussion, the method we present (aligned probing) generalizes to any textual property describing the input and/or output.
2.1 Evaluating Model Behavior
In toxicity research, language model behavior is analyzed via the toxicity of generations. Given the serious implications of toxic language in generations, the standard evaluation protocol considers multiple outputs () for a single input () to capture the model’s worst-case behavior (Gehman et al., 2020; Jain et al., 2024; Gallegos et al., 2024). Following this approach, we generate 25 samples per input using a temperature of and nucleus sampling with (Holtzman et al., 2020). We then evaluate the toxicity of these generations using the PERSPECTIVE API111https://perspectiveapi.com, a widely-used industry standard for toxicity assessment (Wen et al., 2023; Liang et al., 2023; Groeneveld et al., 2024). With these toxicity scores, we compute two metrics:
Expected Maximum Toxicity (EMT)
We compute the maximum toxicity across multiple generations for a given input (). Since captures the model’s worst-case behavior, it answers: How toxic is a language model?
Toxicity Correlation (TC)
We compute the Pearson correlation between the toxicity scores of the input () and the corresponding model toxicity (). This metric quantifies how input toxicity relates to generation toxicity, to answer the question: Do models replicate input toxicity?
2.2 Evaluating Model Internals
To evaluate how models encode information about toxicity, we examine the layer-wise information strength of model internals with respect to toxicity scores. We adopt the probing classifier methodology (Tenney et al., 2019a, b; Belinkov, 2022) and approximate information strength () with the performance of a linear model () that maps internal representations () at each layer to toxicity scores ():
Concretely, we first train222For details on training, see Appendix § A.2. a probe to predict from , where the prediction follows:
We then approximate the encoding strength () as the Pearson correlation between the predicted () and actual () toxicity scores. Since the learning capacity of the probe is limited, a high correlation suggests that substantial information about toxicity is encoded in , while a low correlation indicates weaker encoding. Using this method, we formulate four scenarios (Figure 1) to analyze the encoding of input and output toxicity ( and ) within input and output internals ( and ):
Scenario Input
We first assess how strongly an LM encodes the toxicity of the input within its internals.
Thus, we probe how strongly the input internals () encode information about the input toxicity score (), yielding the information strength .
Scenario Forward
Secondly, we examine how much information about the input’s toxicity is forwarded and retained during generation.
To quantify this, we measure the information strength by probing whether the input toxicity score () is encoded within the internals of the output ().
Scenario Output
The third scenario assesses how much information LMs encode about the toxicity of their generations.
Thus, we measure the information strength by probing whether the output toxicity score () is reflected in the output internals ().
Scenario Backward
Finally, we analyze how much information about output toxicity an LM encodes within its internal representations of the input, i.e.,
we measure the information strength by probing whether the model’s internal representations of the input () strongly encode the aggregated expected maximum toxicity score (EMT) of generations.
2.3 The Interplay of Behavior and Internals
Since we analyze both model behavior and information strength within internal representations for the same toxicity attributes, we can address: How are the internals of models related to their behavior? To quantify this interplay, we examine the relationship between information strength in different probing scenarios () and the model’s toxicity. For instance, if we aim to investigate how strongly the encoding of input toxicity within input internals (Scenario Input) relates to model behavior, we compute the Pearson correlation between the information strength values across models () and their corresponding expected maximum toxicity scores ():
3 Toxic Language
Following Gehman et al. (2020), we define toxic text as text which makes people leave a discussion with high probability. As toxicity is a heterogeneous phenomenon, we focus on six fine-grained attributes: General Toxicity, Identity Attack, Insult, Profanity, Threat, and Sexually Explicit. We quantitatively demonstrate how these attributes capture distinct aspects of toxic language as their score distributions (§ 3.2) and sensitivity to specific tokens (§ 3.3) vary substantially.
Attribute () | Example Prompt | Avg. | |
General Toxicity | I would rather kill them … | ||
Profanity | It’s damn easy to sculpt test … | ||
Insult | Incredibly racist and stupid man … | ||
Sexually Explicit | My butt muscles are so tight … | ||
Identity Attack | If you are white trash … | ||
Threat | If you want to kill … |
3.1 Data
We use the RealToxicPrompts dataset (Gehman et al., 2020) for our analysis and subsequent experiments. This dataset consists of text prompts () paired with corresponding continuations (), each annotated with toxicity scores obtained from the PERSPECTIVE API. We carefully subsample the original 100K samples to optimize computational efficiency while maintaining validity, i.e., we iteratively reduce the dataset size as long as the toxicity scores for all attributes () do not differ statistically significantly () from the full dataset. Following this procedure, our final subset consists of 22K samples.
3.2 Score Distribution
We analyze the score distribution of unique toxicity attributes () within our subset of the RealToxicPrompts dataset. Among all attributes, we find the highest average score for General Toxicity (), suggesting that this attribute is the most sensitive to the PERSPECTIVE API scoring. The average score gradually decreases from Profanity () to Threat, which has the lowest average score (). Additionally, toxicity scores of prompts () and their continuations () marginally correlate, with on average. Thus, the toxicity scores of the prompt and continuation seem unrelated on average, as also shown in Gehman et al. (2020). However, comparing unique toxicity attributes reveals that toxicity scores tend to be replicated within the continuation for Sexually Explicit (), Identity Attack (), and Threat ().
Analyzing the relation among toxicity scores of unique attributes shows strong correlations across General Toxicity, Profanity, and Insult (see Figure 2). In contrast, Threat, Identity Attack, and Sexually Explicit weakly correlate with others. This shows that these scores are complementary and offer a distinct perspective on toxicity.
3.3 Word Sensitivity
We quantify the sensitivity of different toxicity attributes () to individual words. To this end, we retrieve the toxicity scores of a prompt () and separately compute scores for its constituent words . We then define the word sensitivity for a given attribute as the difference between the toxicity score of the prompt () and the toxicity score of its most toxic word:
A high word sensitivity score () indicates that attribute is particularly dependent on individual, presumably explicit, words. Conversely, a low or negative suggests that the attribute captures more contextualized forms of toxic language.
We calculate this word sensitivity for every attribute using all prompts of our dataset. Following Figure 3, General Toxicity, Profanity, and Sexually Explicit are more sensitivity to single word as the average is positive. In contrast, attributes such as Insult, Identity Attack, and Threat have word sensitivity scores centered around zero or negative values, indicating a stronger dependence on the context of a text. The high variance in General Toxicity suggests that it captures a broader spectrum of toxic language, whereas attributes like Sexually Explicit represent more narrowly defined categories. Together with our toxicity score distribution analysis, these insights further highlight the heterogeneous nature of toxic language.
4 Toxicity of Language Models
In this section, we apply aligned probing to comprehensively evaluate LMs in the context of toxicity. We begin by discussing the toxicity of LM generations (§ 4.1), after which we turn to how models encode and propagate information about toxic language internally (§ 4.2). Finally, we connect our behavioral and model-internal insights and study their interplay (§ 4.3).
Setup
We present results aggregated across six popular pre-trained LMs with 7B to 8B parameters from the OLMo, Llama, and Mistral families. See Table 4 in the appendix for more details.
Max. Tox. () | Tox. Corr. (TC) | |||
Attribute | Toxic | Not Toxic | Toxic | Not Toxic |
Average | ||||
General Toxicity | ||||
Profanity | ||||
Insult | ||||
Sexually Explicit | ||||
Identity Attack | ||||
Threat |
4.1 Behavioral Evaluation
We begin by analyzing the toxicity of LMs based on their generated text. Overall, our results (Table 2) align with previous work (Gehman et al., 2020) as LMs generally generate text with substantial toxicity, with EMT of for toxic and for not toxic prompts. Similar to Jain et al. (2024), we find that the input toxicity moderately correlates with the subsequent output toxicity (TC), demonstrating how LMs replicate input properties. Below, we detail our main findings:
i) LMs replicate and amplify toxicity more than human language.
We compare model-generated continuations () with naturally occurring continuations from the RealToxicPrompts dataset () to analyze differences in toxic language between LMs and human language. Our results show that models generate more toxic text than humans do, particularly for toxic prompts, where we observe an increase of in EMT. Furthermore, LM generations replicate input toxicity levels beyond those found in human language. Interestingly, this deviation from human language is similar for both toxic () and not toxic () prompts, suggesting that LMs exhibit fundamentally different behavior from humans, regardless of input toxicity.
ii) LMs are more toxic when single words convey toxicity.
We observe that toxicity levels of LMs vary across the six fine-grained toxicity attributes we consider (Table 2). LMs exhibit particularly high toxicity and strongly replicate input toxicity for attributes sensitive to single words (high in Figure 3). This effect is most pronounced for Sexually Explicit toxic prompts, which show the highest toxicity levels, with EMT and TC scores of and , respectively. In contrast, LMs generate less toxic output and replicate input toxicity to a lesser extent for more context-dependent attributes like Threat and Insult. Additionally, we find that the gap between LMs and human behavior is larger for toxicity that is more explicit (e.g., for Profanity), compared to diffuse attributes like Threat ().
Summary
Our analysis shows that LMs not only replicate but also amplify the toxicity of input prompts, particularly for attributes highly sensitive to single words. This difference among unique types of toxicity demonstrates that LM behavior is as heterogeneous as these attributes themselves.
4.2 Internal Evaluation
Now that we have analyzed LM behavior, we turn to how they encode toxic language internally.
iii) Toxic language is encoded in lower layers.
Figure 4 illustrates how strongly (lines) and consistently (bands) LMs encode the toxicity of text based on the average and standard deviation across 20 probes covering multiple folds and seeds. Our findings challenge previous research that attributes toxicity encoding to upper LM layers (Lee et al., 2024). Instead, we observe three-stages: (1) information emerges and peaks in the first third of model layers, (2) gradually declines in the middle third, and (3) continues decreasing in later layers while standard deviation increases. Notably, the standard deviation (bands) reveals differences even in layers with similar information strength, such as layer one and layer 23, which exhibit deviations of and , respectively. These variations suggest inherent differences across model regions and highlight the necessity of thorough evaluations. We validate these insights with alternative probing metrics, namely selectivity (Hewitt and Liang, 2019) and compression (Voita and Titov, 2020) - see Figure 17 and Figure 16 in the appendix.
iv) Information strength varies by toxicity attribute.
We further analyze how the encoding of toxic language differs across specific toxicity attributes. As shown in Figure 5, LMs encode less information for contextualized attributes, such as Threat, while attributes with higher word sensitivity, like General Toxicity, are more strongly encoded. This observation aligns with prior work (Warstadt et al., 2020; Waldis et al., 2024b), which found that LMs encode word-level properties, such as morphology, more strongly than contextual information. Interestingly, the maximum information strength for contextualized attributes occurs in higher layers, such as layer 7 for Identity Attack. In contrast, attributes sensitive to single words, such as Sexually Explicit, peak in lower layers, which probably capture such surface features (Tenney et al., 2019a; Niu et al., 2022).
v) LMs know more about input toxicity and propagate this information.
Our analysis (Figure 4) shows that LMs encode more information about input toxicity () than output toxicity (). This information strength reaches up to in the Input scenario and in Forward, while it is lower for output toxicity, with a maximum of in Output and in Backward. These findings build on previous work (West et al., 2024) and suggest that LMs struggle to internalize the meaning of their outputs to toxicity.
At the same time, our results show that LMs not only encode input toxicity strongly in input internals () but also transfer this information to generation internals (). This is particularly clear when comparing the Forward and Output scenarios, where input toxicity () is encoded almost as strongly as output toxicity () in the output internals. Additionally, the delayed rise of information in output internals supports this transfer: it takes six layers to exceed an information strength of in the Forward scenario, indicating that LMs gradually pass this information through the attention mechanism. This confirms that LMs entangle their generations with input toxicity, emphasizing the need to understand better how toxicity is encoded and transferred within models.
Summary
Our insights reveal that model internals strongly encode toxic language, especially input toxicity, and attribute sensitivity to single words. Additionally, model layers vary in information strength, clarity, and the encoding of unique toxicity attributes.
4.3 Correlation of Internals and Behavior
Connecting our behavioral and internal evaluations, we show that information strength is closely related to observable toxicity when comparing distinct toxicity attributes. Figure 6 demonstrates that model toxicity for specific attributes () increases when their internals (, ) encode more information about . This correlation is stronger in the Input, Forward, and Output scenarios, reaching up to , , and , respectively, while it is lower for Backward (). These findings suggest that encoding input and output toxicity for a specific attribute () more strongly increases the model toxicity related to .
5 Comparing Language Models
After evaluating toxicity in general, we next examine how individual models differ. In § 5.1, we discuss insights about how the behavior of specific LMs varies, with a particular focus on the effects of instruction tuning. We then present findings on how internals differ (§ 5.2) and finally analyze the interplay between model internals and behavior across distinct LMs (§ 5.3).
Setup
We evaluate pre-trained and instruction-tuned versions of the following popular contemporary models: OLMo, OLMo-2, Llama-2, Llama-3, Llama-3.1, and Mistral-v0.3.333See Table 4 of the appendix for details. With each model, we discuss results averaged across the six fine-grained toxicity attributes.
5.1 Behavioral Evaluation
We first analyze how the behavior of unique models differs in the context of toxicity, with a focus on how instruction-tuning changes LMs.
i) Instruction-tuning diversifies LMs.
Comparing LMs reveals only minor differences in toxicity among pre-trained LMs (see Table 6 of the appendix). Notably, OLMo exhibits the lowest toxicity, highlighting the effectiveness of carefully curated, detoxified pre-training data (Groeneveld et al., 2024). In contrast, instruction-tuned LMs show more behavioral variation, especially for toxic prompts. These differences are particularly pronounced for LMs presumably trained on distinct instruction corpora, such as Llama-2-Chat and OLMo-Instruct. As these results underline the impact of pre-training and instruction-tuning data, only releasing these corpora would allow us to examine LMs and their limitations holistically.
ii) Instruction-tuning mitigates toxicity.
Consistent with Jain et al. (2024), instruction-tuned (IT) LMs exhibit lower toxicity than pre-trained (PT) ones, with an of for toxic prompts and for not toxic prompts (see Table 3). In fact, the toxicity of IT LMs is more closely aligned with the toxicity of human language for toxic prompts () while being lower () for not toxic prompts. Analyzing the correlation with input toxicity (TC) reveals that IT models effectively suppress high input toxicity ( for toxic prompts) while preserving the low toxicity of not toxic prompts ().
Since IT LMs frequently generate phrases like as a helpful assistant, this mitigation effect may partly stem from such formulations. Re-evaluating generations without such phrases results in a slight increase in toxicity (see Figure 10 in the appendix). However, their toxicity remains lower than pre-trained LMs, demonstrating that instruction-tuning reduces LM toxicity without explicit objectives beyond exposure to presumably not toxic preference data. Interestingly, this adaptation appears more implicit, as toxicity mitigation is particularly pronounced for more contextually nuanced attributes such as Threat.
Max. Tox. (EMT) | Tox. Corr. (TC) | |||
Language Model | Toxic | Not Toxic | Toxic | Not Toxic |
Avg. Pre-Trained (PT) | ||||
Avg. Instruction-Tuned (IT) |
Summary
These insights show that instruction-tuning effectively mitigates toxic language, and this subsequent stage, after pre-training, shapes behavioral differences across unique models.
5.2 Internal Evaluation
Next, we analyze how LMs encode toxic language differently, grouped by whether they are just pre-trained or also instruction-tuned.
iii) LMs differ in how they encode toxicity in upper layers.
Analyzing how LMs encode toxic language, we find that they exhibit similar encoding patterns in lower layers but diverge in upper layers (Figure 7). Notably, as this pattern holds for both pre-trained (PT) and instruction-tuned (IT) models, it contrasts with the behavioral similarities across PT models. We assume these upper layers encode more information about output semantics, potentially resulting in similar toxicity scores. Moreover, this finding aligns with our previous finding that regions within LMs differ substantially (§ 4.2).
Focusing on individual LMs reveals further model-specific insights. Llama-2 encodes toxicity less strongly and with higher variability than Llama-3 and Llama-3.1, likely due to its smaller pre-training dataset (2T vs. 15T+ tokens). Meanwhile, OLMo exhibits high information strength and low variance, another sign of the high quality of its pre-training data.
iv) Instruction-tuned LMs encode more information about input toxicity.
We compare PT and IT LMs to assess the impact of instruction-tuning on model internals. As shown in Figure 7, instruction-tuning increases the information strength for input toxicity while reducing it for output toxicity, particularly in the Forward and Backward scenarios and in upper layers. Interestingly, the difference between PT and IT LMs is stronger for toxicity attributes that are less sensitive to individual words, especially Threat and Insult. These findings suggest that instruction-tuning primarily affects upper layers, which encode broader linguistic context, rather than lower layers, which focus more on lexical features.
Summary
We find that individual LMs encode information about toxic language more differently from each other in upper layers, while showing more similarity in lower layers. This variance is particularly evident after instruction-tuning, which adapts LMs to encode more information about the input and less about the output toxicity.
5.3 Interplay of Internals and Behavior.
Finally, we correlate the average information strength at each layer with the resulting output toxicity (EMT) across different LMs. As shown in Figure 8, less toxic LMs tend to encode more information about input toxicity, particularly in the Forward scenario and for toxic prompts (). Conversely, these less toxic LMs encode less information about output toxicity, especially in the Backward scenario, where we observe for toxic prompts. These findings suggest that models are generally less toxic when they know more about input toxicity, particularly for attributes with higher word sensitivity, such as Sexually Explicit or Profanity.
6 Correlation or Causation?
So far, we have seen that LMs propagate toxicity from their inputs to their outputs, and their internals strongly correlate with observable toxicity. To establish whether this connection between the internal and behavioral perspectives is causal, we perform layer-wise interventions. Specifically, we measure model toxicity when skipping one layer at a time, approximating the impact of information encoded at that layer. As these experiments are computationally expensive, we focus on the pre-trained OLMo model and focus on layers to , which encode toxic language particularly strongly.
As Figure 9 shows, removing information by skipping model layers generally increases the toxicity of generated text. Specifically, we observe an average increase of in maximum expected toxicity (EMT) across all intervened layers, with a peak of for layer . Relating this to our internal analysis, layer strongly encodes input toxicity in both the input and output. Comparing the results for different toxicity attributes, we confirm that the interplay between model internals and behavior varies across distinct attributes. As shown in § 5.3, this interplay is stronger for explicit attributes, where we observe a more pronounced causal effect. Specifically, removing information causes up to more toxicity for Profanity. In contrast, more contextualized attributes, such as Threat, exhibit only a minor increase.
These findings extend previous insights and suggest information about input toxicity causally enables language models to generate less toxic text. At the same time, these insights underscore the importance of studying causal mechanisms of LMs (Saphra and Wiegreffe, 2024), particularly for safety aspects (Bereska and Gavves, 2024), as LMs vary in how they process distinct toxicity attributes.
7 Case Studies
Finally, we present four case studies with practical applications of aligned probing, focusing on DPO-based detoxification (§ 7.1), multi-prompt evaluation (§ 7.2), model quantization (§ 7.3), and pre-training dynamics (§ 7.4).
7.1 Case Study: Detoxification
We study how model internals change under DPO detoxification in Figure 12 of the appendix. Our results confirm this method’s effectiveness in reducing the toxicity of LMs (Li et al., 2024). However, we find a substantial information loss within the internals of these models, particularly in the upper layers. As we observe this information loss for text properties other than toxicity, like input length, we see that detoxification via DPO impacts model internals substantially. Therefore, a more holistic evaluation is indispensable to quantify what abilities alignment methods remove from models. As such, aligned models can also easily be unaligned (Lee et al., 2024).
7.2 Case Study: Multi-Prompt Evaluation
We study how multi-prompt evaluation impacts the internals and behavior of models by prompting LMs to complete a given text chunk with four different prompt formulations – see Figure 13 of the appendix. These experiments show that the toxicity of LMs varies across different prompts, while model internals remain more stable. These results expand previous work about the crucial entanglement of model behavior and specific instructions (Mizrahi et al., 2024; Sclar et al., 2024). Our results show that this variance is visible beyond task-specific evaluation, and that, in contrast, model internals reveal fewer deviations.
7.3 Case Study: Model Quantization
We also study whether evaluating model internals and behavior vary when we apply quantization methods to improve efficiency – see Figure 14 of the appendix. We find that both behavioral and internal results remain valid and consistent, as we found only minor deviations when comparing full precision with half and four bit precision.
7.4 Case Study: Pre-Training Dynamics
We analyze how model behavior and internals evolve during pre-training by studying six pre-training checkpoints of OLMo (Groeneveld et al., 2024) - see Figure 12 of the appendix. These results show that early in training (100K steps), models are close to their final toxicity and information strength regarding toxic language. Afterward, we mainly see improvements in the clarity of the information strength, with lower standard deviations across folds and seeds after 100K steps. These observations suggest that aligned probing can effectively monitor pre-training dynamics.
8 Related Work
Toxicity of Language Models
Work on language model toxicity primarily focuses on evaluating and modifying model behavior by analyzing inputs and outputs (Gallegos et al., 2024). For instance, Gehman et al. (2020) examine toxicity in generations given English prompts, while de Wynter et al. (2024) and Jain et al. (2024) extend this to multilingual settings. Wen et al. (2023) go beyond overt toxicity, investigating implicit toxicity that is harder for automatic classifiers to detect. Another line of research explores the origins of toxicity in LMs by analyzing training data. Gehman et al. (2020) highlight the prevalence of toxic content in pre-training corpora, and Longpre et al. (2024) show that filtering for quality and toxicity can paradoxically lead to toxic degeneration and poor generalization. Unlike these works, we comprehensively evaluate LMs by relating the study of their behavior and model internals, with different types of toxic language.
Studying Model Internals
Recent interpretability research has begun probing toxicity within model internals. Ousidhoum et al. (2021) first explored this by using masked language models. More recent work analyzes and mitigates toxicity via model merging (Yang et al., 2024), direct preference optimization (DPO) (Lee et al., 2024; Li et al., 2024), and knowledge editing (Wang et al., 2024). Methods such as linear probing, activation analysis, and causal interventions have been used to study toxicity mitigation in both English (Lee et al., 2024) and multilingual models (Li et al., 2024). While we adopt similar methods, we contribute a new framework, aligned probing, to trace toxicity through model internals, enabling a deeper understanding of how input toxicity is entangled with subsequent model behavior.
Probing
Our approach builds on classifier-based probing, which has been widely studied (Belinkov, 2022). Probing classifiers can be difficult to interpret, leading to refinements such as control tasks (Hewitt and Liang, 2019; Ravichander et al., 2021), fine-tuning probes (Mosbach et al., 2020), information-theoretic perspectives (Voita and Titov, 2020), and behavioral explanations (Elazar et al., 2021). While our study focuses on toxicity, probing has been applied to various linguistic properties, including negation and function words (Kim et al., 2019), grammatical number (Lasri et al., 2022), author demographics (Lauscher et al., 2022), language identity (Srinivasan et al., 2023), topic classification (Waldis et al., 2024a), and linguistic competence (Waldis et al., 2024b).
9 Discussion and Conclusion
We present aligned probing, a method to trace text properties from the model input to the output and connect these findings to subsequent behavior. By applying this method in the context of toxicity, we evaluate over 20 contemporary models and demonstrate that they substantially encode information about toxic language, which crucially impacts the toxicity of model outputs. Moreover, our results reveal that model behavior strongly relies on the toxicity of the input, and model internals strongly encode and propagate information about this input toxicity. With this substantial dependence on the properties of the input text, we identify a crucial dilemma of generative models: We expect them to generate a semantically relevant output given an input prompt without considering unwanted properties, such as toxicity. Pursuing this thought towards more controllable text generation, we plan to apply aligned probing to analyze other aspects of generation, like stereotypical formulations, and examine the nature of other mitigation methods, such as model merging.
Acknowledgements
We thank Pia Pachinger, Ilia Kuznetsov, Tim Baumgärtner, and Paul Röttger for their valuable feedback and discussions. Andreas Waldis is supported by the Hasler Foundation Grant No. 21024. The work of Anne Lauscher is funded under the Excellence Strategy of the German Federal Government and States.
Appendix A Appendix
A.1 Limitations
Classifying Toxicity
Detecting toxicity is a non-trivial task as conceptualizations, datasets, and annotator attitudes can vary widely (Waseem, 2016; Waseem et al., 2017; Sap et al., 2022; Pachinger et al., 2023; Cercas Curry et al., 2024). Moreover, toxicity - as with most linguistic properties - is highly contextual and can be implicit, making it difficult to detect (Wen et al., 2023). Even though we consider fine-grained toxicity attributes, our use of PERSPECTIVE API444An industry standard API providing high performance in toxicity detection, see results online. and probing classifiers may miss forms of toxicity not represented in upstream datasets, or exhibit biases (Nogara et al., 2023; Pozzobon et al., 2023).
Beyond English Toxicity
Due to space constraints, we only demonstrate aligned probing with English toxicity in this paper, but the framework is broad; it can be applied to any language model and any textual property. We emphasize that English toxicity is intended as an example, and other English textual properties may be encoded and propagated differently from input to output through model internals. Additionally, evaluations of toxicity in non-English languages are also influenced by whether the data is localized to linguistically and culturally appropriate examples and can still be affected by English pre-training data (Jain et al., 2024).
Probing Classifiers
One of the fundamental problems with using probing classifiers is their limited utility for model explainability. In other words, just because a model’s representations are predictive of a property does not mean that the model is using it (Ravichander et al., 2021; Elazar et al., 2021; Belinkov, 2022). We address this limitation by correlating probing performance with actual toxic behavior when presenting our results, by running causal analyses in addition to correlative analyses, by using control tasks (Hewitt and Liang, 2019) and evaluating our probing setup from an information theory perspective (Voita and Titov, 2020). In future applications of aligned probing to other text properties, it is important to contextualize results with these checks as we do.
A.2 Experimental Details
Probing Hyperparameters
We use fixed hyperparameters for training the probes following previous work (Hewitt and Liang, 2019; Voita and Titov, 2020). Specifically, we train for 20 epochs, selecting the optimal one based on development instances. We use AdamW (Loshchilov and Hutter, 2019) as the optimizer, with a batch size of 16, a learning rate of 0.001, a dropout rate of 0.2, and a warmup phase covering 10% of the total steps. Additionally, we set the random seeds to .
Model | Huggingface Tag | Parameters | Pre-Training Tokens |
OLMo-5k (Groeneveld et al., 2024) | allenai/OLMo-7B-hf | 7 billion | 0.35T tokens |
OLMo-100k (Groeneveld et al., 2024) | allenai/OLMo-7B-hf | 7 billion | 0.7T tokens |
OLMo-200k (Groeneveld et al., 2024) | allenai/OLMo-7B-hf | 7 billion | 1.05T tokens |
OLMo-300k (Groeneveld et al., 2024) | allenai/OLMo-7B-hf | 7 billion | 1.4T tokens |
OLMo-400k (Groeneveld et al., 2024) | allenai/OLMo-7B-hf | 7 billion | 1.75T tokens |
OLMo-500k (Groeneveld et al., 2024) | allenai/OLMo-7B-hf | 7 billion | 2.1T tokens |
OLMo (Groeneveld et al., 2024) | allenai/OLMo-7B-hf | 7 billion | 2.5T tokens |
OLMo-Instruct (Groeneveld et al., 2024) | allenai/OLMo-7B-Instruct-hf | 7 billion | 2.5T tokens + 381k instructions |
[…]
OLMo-2-Instruct (OLMo et al., 2025) | allenai/OLMo-2-1124-7B-Instruct | 7 billion | 4.1T tokens + 367k instructions |
[…]
Llama-2-Chat (Touvron et al., 2023) | meta-llama/Llama-2-7b-chat-hf | 7 billion | 2T tokens + 1.4M instructions |
Llama-2-Detox (Rafailov et al., 2023) | BatsResearch/llama2-7b-detox-qlora | 7 billion | 2T tokens + 25k demonstrations |
[…]
Llama-3-Instruct (Grattafiori et al., 2024) | meta-llama/Meta-Llama-3-8B-Instruct | 8 billion | 15T+ tokens + unknown instructions |
Llama-3-Detox (Rafailov et al., 2023) | BatsResearch/llama3-8b-detox-qlora | 8 billion | 15T+ tokens + 25k demonstrations |
[…]
Llama-3.1-Instruct (Grattafiori et al., 2024) | meta-llama/Llama-3.1-8B-Instruct | 8 billion | 15+ tokens + unknown instructions |
[…]
Mistral-v0.3-Instruct (Jiang et al., 2023) | mistralai/Mistral-7B-Instruct-v0.3 | 7 billion | unknown |
[…]
Aya-23-Detox (Rafailov et al., 2023) | BatsResearch/aya-23-8b-detox-qlora | 8 billion | unknown + 5.7M instructions + 25k demonstrations |
Hardware
All experiments are conducted on 20 Nvidia RTX A6000 GPUs. Each GPU is equipped with 48GB of memory and 10,752 CUDA cores.
Considered LMs
Table 4 provides an overview of the language models considered in this study.
Your task is to complete the given text without explanation. Text: {Text} |
Complete the following text without explanation. Text: {Text} |
Complete a given text without explanation. Text: {Text} |
Complete the following text without explanation. Text: {Text} |
Max. Tox. (EMT) | Tox. Corr. (TC) | |||
Attribute () | Toxic | Not Toxic | Toxic | Not Toxic |
OLMo | ||||
OLMo-Instruct | ||||
[…]
OLMo-2-Instruct | ||||
[…]
Llama-2-Chat | ||||
[…]
Llama-3-Instruct | ||||
[…]
Llama-3.1-Instruct | ||||
[…]
Mistral-v0.3-Instruct |
Case Study 1: Detoxification
a. Behavioral Results
Max. Tox. (EMT) | Tox. Corr. (TC) | |||
Attribute () | Toxic | Not Toxic | Toxic | Not Toxic |
Llama-2 | ||||
Llama-2-Chat | ||||
Llama-2-Detox | ||||
[…]
Llama-3-Instruct | ||||
Llama-3-Detox | ||||
[…]
Aya-23-Detox |
b. Internal Results
Case Study 2: Multi-Prompt Evaluation
a. Behavioral Results
Identity Attack | Insult | Profanity | Sexually Explicit | Threat | General Toxicity | |||||||
Model | Toxic | Not Toxic | Toxic | Not Toxic | Toxic | Not Toxic | Toxic | Not Toxic | Toxic | Not Toxic | Toxic | Not Toxic |
OLMo-Instruct | ||||||||||||
OLMo-2-Instruct | ||||||||||||
Llama-2-chat | ||||||||||||
Llama-3-Instruct | ||||||||||||
Llama-3.1-Instruct | ||||||||||||
Mistral-v0.3-Instruct |
b. Internal Results
Case Study 3: Model Quantization
a. Behavioral Results
Identity Attack | Insult | Profanity | Sexually Explicit | Threat | General Toxicity | |||||||
Model | Toxic | Not Toxic | Toxic | Not Toxic | Toxic | Not Toxic | Toxic | Not Toxic | Toxic | Not Toxic | Toxic | Not Toxic |
OLMo-Full | ||||||||||||
OLMo-Half | ||||||||||||
OLMo-Four-Bit |
b. Internal Results
Case Study 4: Pre-Training Dynamics
a. Behavioral Results
Identity Attack | Insult | Profanity | Sexually Explicit | Threat | General Toxicity | |||||||
Model | Toxic | Not Toxic | Toxic | Not Toxic | Toxic | Not Toxic | Toxic | Not Toxic | Toxic | Not Toxic | Toxic | Not Toxic |
OLMo-5k | ||||||||||||
OLMo-100k | ||||||||||||
OLMo-200k | ||||||||||||
OLMo-300k | ||||||||||||
OLMo-400k | ||||||||||||
OLMo-500k | ||||||||||||
OLMo-Full |
b. Internal Results

### v2

@@
 How do LMs encode information about toxicity, and what is the interplay between their internals and behavior?
-We address this gap by introducing aligned probing (Figure 1), a novel interpretability framework (§ 2) that aligns model behavior with internal representations for toxicity.
+We address this gap by introducing aligned probing (Figure 1), a novel interpretability framework (§ 2) that aligns model behavior with internals for toxicity.
 First, we prompt LMs with inputs and assess the toxicity of their generated outputs.
-Then, during the forward pass, we extract internal representations at each layer to analyze how models encode toxic language.
-Specifically, we use linear probing (Tenney et al., 2019a; Belinkov, 2022) to train linear models to use these internals to predict specific properties (like input toxicity).
-Since probes have limited capacity and are rigorously validated (Hewitt and Liang, 2019; Voita and Titov, 2020), their prediction performance on held-out data estimates information strength.
+During this forward pass, we extract internal representations to analyze how models encode toxic language.
+Specifically, we extract the hidden states at each model layer and average them for all input and output tokens.
+Then, we use linear probing (Tenney et al., 2019a; Belinkov, 2022) to estimate the encoded information about the toxicity of the input or output.
+As we train linear models (probes) with limited capacity and rigorously validate them (Hewitt and Liang, 2019; Voita and Titov, 2020), their ability to estimate toxicity based on internal effects effectively approximates information strength.
 Finally, we relate the behavioral and internal perspectives, examining their interplay.
@@
 2.2 Evaluating Model Internals
-To evaluate how models encode information about toxicity, we examine the layer-wise information strength of model internals with respect to toxicity scores.
-We adopt the probing classifier methodology (Tenney et al., 2019a, b; Belinkov, 2022) and approximate information strength () with the performance of a linear model () that maps internal representations () at each layer to toxicity scores ():
-Concretely, we first train222For details on training, see Appendix § A.2. a probe to predict from , where the prediction follows:
+To evaluate how models encode information about toxicity, we extract the hidden states () averaged across all input () and output tokens () at every model layer .
+We then adopt the probing classifier methodology (Tenney et al., 2019a, b; Belinkov, 2022) to assess how these these internals () linearly map to the corresponding toxicity scores ():
+Concretely, we first train222For details on training, see Appendix § A.1. a probe (linear model) to predict from , where the prediction follows:
 We then approximate the encoding strength () as the Pearson correlation between the predicted () and actual () toxicity scores.
 Since the learning capacity of the probe is limited, a high correlation suggests that substantial information about toxicity is encoded in , while a low correlation indicates weaker encoding.
-Using this method, we formulate four scenarios (Figure 1) to analyze the encoding of input and output toxicity ( and ) within input and output internals ( and ):
+Using this method, we formulate four scenarios (Figure 1) to analyze the encoding of input and output toxicity ( and ) within the averaged input and output internals ( and ):
 Scenario Input
@@
 | Identity Attack | ||||
-Threat |
+| Threat | ||||
 4.1 Behavioral Evaluation
@@
 Figure 4 illustrates how strongly (lines) and consistently (bands) LMs encode the toxicity of text based on the average and standard deviation across 20 probes covering multiple folds and seeds.
-Our findings challenge previous research that attributes toxicity encoding to upper LM layers (Lee et al., 2024).
-Instead, we observe three-stages: (1) information emerges and peaks in the first third of model layers, (2) gradually declines in the middle third, and (3) continues decreasing in later layers while standard deviation increases.
+We observe three-stages: (1) information emerges and peaks in the first third of model layers, (2) gradually declines in the middle third, and (3) continues decreasing in later layers while standard deviation increases.
 Notably, the standard deviation (bands) reveals differences even in layers with similar information strength, such as layer one and layer 23, which exhibit deviations of and , respectively.
-These variations suggest inherent differences across model regions and highlight the necessity of thorough evaluations.
 We validate these insights with alternative probing metrics, namely selectivity (Hewitt and Liang, 2019) and compression (Voita and Titov, 2020) - see Figure 17 and Figure 16 in the appendix.
+With these findings, we offer an alternative perspective on toxicity in LMs, one that differs from localizing it in the upper layer by projecting hidden states to specific output vocabulary (Lee et al., 2024).
+Instead, our results suggest that lower model layers compute rich latent information while the upper layers then project this information to text, naturally affected by information diminishing due to the large discrete output space (vocabulary).
+Since the meaning attached to these tokens does not necessarily represent internal information (Shojaee et al., 2025; Hewitt et al., 2025; Kambhampati et al., 2025), supplementary evaluations focusing on latent encoded information are indispensable.
 iv) Information strength varies by toxicity attribute.
@@
 | Avg. Pre-Trained (PT) | ||||
-Avg. Instruction-Tuned (IT) |
+| Avg. Instruction-Tuned (IT) | ||||
 Summary
@@
 Pursuing this thought towards more controllable text generation, we plan to apply aligned probing to analyze other aspects of generation, like stereotypical formulations, and examine the nature of other mitigation methods, such as model merging.
+Limitations
+Classifying Toxicity
+Detecting toxicity is a non-trivial task as conceptualizations, datasets, and annotator attitudes can vary widely (Waseem, 2016; Waseem et al., 2017; Sap et al., 2022; Pachinger et al., 2023; Cercas Curry et al., 2024).
+Moreover, toxicity - as with most linguistic properties - is highly contextual and can be implicit, making it challenging to detect (Wen et al., 2023).
+Even though we consider fine-grained toxicity attributes, our use of PERSPECTIVE API444An industry standard API providing high performance in toxicity detection, see results online.
+and probing classifiers may miss forms of toxicity not represented in upstream datasets, or exhibit biases (Nogara et al., 2023; Pozzobon et al., 2023).
+Probing Classifiers
+This work relies on probing classifiers to assess the information encoded within internal representations, but this comes with methodological constraints.
+First, we want the probe to catch relevant patterns in internal representations of LMs, but it should not learn nonexistent patterns due to its learning capabilities.
+To address this, we use the simplest possible probe – a linear model without intermediate layers – and verify the low learning capabilities with control tasks (Hewitt and Liang, 2019).
+Moreover, we validate our probing setup from an information-theoretic perspective (Voita and Titov, 2020), thereby confirming our results and findings.
+Second, probing is a solely observational explainability method which does not offer any insights about causal relations.
+In other words, just because a model’s representations are predictive of a property does not mean that the model is using it (Ravichander et al., 2021; Elazar et al., 2021; Belinkov, 2022).
+We address this limitation by correlating probing performance with actual toxic behavior when presenting our results, and by running interventions to analyze whether correlations are causal.
+In future applications of aligned probing to other text properties, it is important to contextualize results with these checks as we do.
+Beyond English Toxicity
+Due to space constraints, we only demonstrate aligned probing with English toxicity in this paper.
+We emphasize that English toxicity is intended as an example, and other English textual properties may be encoded and propagated differently from input to output through model internals.
+Additionally, evaluations of toxicity in non-English languages are also influenced by whether the data is localized to linguistically and culturally appropriate examples and can still be affected by English pre-training data (Jain et al., 2024).
 Acknowledgements
@@
 Appendix A Appendix
-A.1 Limitations
-Classifying Toxicity
-Detecting toxicity is a non-trivial task as conceptualizations, datasets, and annotator attitudes can vary widely (Waseem, 2016; Waseem et al., 2017; Sap et al., 2022; Pachinger et al., 2023; Cercas Curry et al., 2024).
-Moreover, toxicity - as with most linguistic properties - is highly contextual and can be implicit, making it difficult to detect (Wen et al., 2023).
-Even though we consider fine-grained toxicity attributes, our use of PERSPECTIVE API444An industry standard API providing high performance in toxicity detection, see results online.
-and probing classifiers may miss forms of toxicity not represented in upstream datasets, or exhibit biases (Nogara et al., 2023; Pozzobon et al., 2023).
-Beyond English Toxicity
-Due to space constraints, we only demonstrate aligned probing with English toxicity in this paper, but the framework is broad; it can be applied to any language model and any textual property.
-We emphasize that English toxicity is intended as an example, and other English textual properties may be encoded and propagated differently from input to output through model internals.
-Additionally, evaluations of toxicity in non-English languages are also influenced by whether the data is localized to linguistically and culturally appropriate examples and can still be affected by English pre-training data (Jain et al., 2024).
-Probing Classifiers
-One of the fundamental problems with using probing classifiers is their limited utility for model explainability.
-In other words, just because a model’s representations are predictive of a property does not mean that the model is using it (Ravichander et al., 2021; Elazar et al., 2021; Belinkov, 2022).
-We address this limitation by correlating probing performance with actual toxic behavior when presenting our results, by running causal analyses in addition to correlative analyses, by using control tasks (Hewitt and Liang, 2019) and evaluating our probing setup from an information theory perspective (Voita and Titov, 2020).
-In future applications of aligned probing to other text properties, it is important to contextualize results with these checks as we do.
-A.2 Experimental Details
+A.1 Experimental Details
 Probing Hyperparameters
@@
 | OLMo-Instruct (Groeneveld et al., 2024) | allenai/OLMo-7B-Instruct-hf | 7 billion | 2.5T tokens + 381k instructions |
-[…]
+| OLMo-2 (OLMo et al., 2025) | allenai/OLMo-2-1124-7B | 7 billion | 4.1T tokens |
 | OLMo-2-Instruct (OLMo et al., 2025) | allenai/OLMo-2-1124-7B-Instruct | 7 billion | 4.1T tokens + 367k instructions |
-[…]
+| Llama-2 (Touvron et al., 2023) | meta-llama/Llama-2-7b-hf | 7 billion | 2T tokens |
 | Llama-2-Chat (Touvron et al., 2023) | meta-llama/Llama-2-7b-chat-hf | 7 billion | 2T tokens + 1.4M instructions |
 | Llama-2-Detox (Rafailov et al., 2023) | BatsResearch/llama2-7b-detox-qlora | 7 billion | 2T tokens + 25k demonstrations |
-[…]
+| Llama-3 (Grattafiori et al., 2024) | meta-llama/Meta-Llama-3-8B-Instruct | 8 billion | 15T+ tokens |
 | Llama-3-Instruct (Grattafiori et al., 2024) | meta-llama/Meta-Llama-3-8B-Instruct | 8 billion | 15T+ tokens + unknown instructions |
 | Llama-3-Detox (Rafailov et al., 2023) | BatsResearch/llama3-8b-detox-qlora | 8 billion | 15T+ tokens + 25k demonstrations |
-[…]
+| Llama-3.1 (Grattafiori et al., 2024) | meta-llama/Llama-3.1-8B | 8 billion | 15T+ tokens |
 | Llama-3.1-Instruct (Grattafiori et al., 2024) | meta-llama/Llama-3.1-8B-Instruct | 8 billion | 15+ tokens + unknown instructions |
-[…]
+| Mistral-v0.3 (Jiang et al., 2023) | mistralai/Mistral-7B-v0.3 | 7 billion | unknown |
 | Mistral-v0.3-Instruct (Jiang et al., 2023) | mistralai/Mistral-7B-Instruct-v0.3 | 7 billion | unknown |
-[…]
+| Aya-23 (Aryabumi et al., 2024) | CohereForAI/aya-23-8B | 8 billion | unknown + 5.7M instructions |
 | Aya-23-Detox (Rafailov et al., 2023) | BatsResearch/aya-23-8b-detox-qlora | 8 billion | unknown + 5.7M instructions + 25k demonstrations |
@@
 | OLMo-Instruct | ||||
-[…]
+| OLMo-2 | ||||
 | OLMo-2-Instruct | ||||
-[…]
+| Llama-2 | ||||
 | Llama-2-Chat | ||||
-[…]
+| Llama-3 | ||||
 | Llama-3-Instruct | ||||
-[…]
+| Llama-3.1 | ||||
 | Llama-3.1-Instruct | ||||
-[…]
-Mistral-v0.3-Instruct |
+| Mistral-v0.3 | ||||
+| Mistral-v0.3-Instruct | ||||
 Case Study 1: Detoxification
@@
 | Llama-2-Detox | ||||
-[…]
+| Llama-3 | ||||
 | Llama-3-Instruct | ||||
 | Llama-3-Detox | ||||
-[…]
-Aya-23-Detox |
+| Aya-23 | ||||
+| Aya-23-Detox | ||||
 b. Internal Results
@@
 | Llama-3.1-Instruct | ||||||||||||
-Mistral-v0.3-Instruct |
+| Mistral-v0.3-Instruct | ||||||||||||
 b. Internal Results
@@
 | OLMo-Half | ||||||||||||
-OLMo-Four-Bit |
+| OLMo-Four-Bit | ||||||||||||
 b. Internal Results
@@
 | OLMo-500k | ||||||||||||
-OLMo-Full |
+| OLMo-Full | ||||||||||||
 b. Internal Results